2. There were 1500 people at a high school football game. Student tickets were $2.00 and adult tickets were $3.50. The total receipts for the game were $3825. How many students bought tickets?

Let a = the number of adult tickets purchased.
Let s = the number of student tickets purchased.
a. Write a system of equations that can be used to determine the number of adult and student tickets purchased.
b. Determine the number of adults tickets sold and the number of student tickets sold. Use mathematics to explain how you determined your answer.


Answer by checkley71(8403)

A+S=1500 OR A=1500-S
3.5A+2S=3825
NOW REPLACE A WITH(1500-S) & SOLVE FOR S.
3.5(1500-S)+2S=3825
5250-3.5S+2S=3825
-1.5S=3825-5250
-1.5S=1425
S=-1425/-1.5
S=950 IS THE NUMBER OF CHILDREN TICKETS SOLD.
1500-950=550 NUMBER OF ADULT TICKETS SOLD.
PROOF.
3.5*550+2*950=3825
1925+1900=3825
3825=3825